In a recent school year in the state of Washington, there were 325,000 high school students. Of these, 156,000 were girls and 169,000 were boys. Among the girls, 54,500 dropped out of school, and among the boys, 10,400 dropped out. A student is chosen at random. Round the answers to four decimal places. (a) What is the probability that the student is male? (b) What is the probability that the student dropped out? (c) What is the probability that the student is male and dropped out? (d) Given that the student is male, what is the probability that he dropped out? (e) Given that the student dropped out, what is the probability that the student is male?
Answer #1

(a) the probability that the student is male = 169000/325000 =0.52

(b) the probability that the student dropped out = (54500+10400) / 325000 = 0.1997

(c) the probability that the student is male and dropped out = 10400/325000 = 0.032

(d) Given that the student is male, the probability that he dropped out = 10400/169000 = 0.6154

(e) Given that the student dropped out, the probability that the student is male = 10400/(10400+54500) = 0.1602